OpenStudy (anonymous):

Find the derivative y' of the function de fined implicitly by:

OpenStudy (anonymous):

\[\sqrt{x+y}+\sqrt{xy}=6\] now this is what i have so far: \[(\frac{ 1 }{ 2 }(x+y)^\frac{ -1 }{ 2 }*y'+\frac{ 1 }{ 2\sqrt{xy} }(1*y+y'*x)=0\] so then i get: \[\frac{ (x+y)^\frac{ -1 }{ 2 }*y'*\sqrt{xy} }{ 2\sqrt{xy} }+\frac{ y }{ 2\sqrt{xy} }+\frac{ xy' }{ 2\sqrt{xy} }=0\]

OpenStudy (anonymous):

not sure if i have the third step done correctly, and not sure where to go from here.

hartnn (hartnn):

(x+y)' = (1+y')

hartnn (hartnn):

\((\frac{ 1 }{ 2 }(x+y)^\frac{ -1 }{ 2 }*(1+y')+\frac{ 1 }{ 2\sqrt{xy} }(1*y+y'*x)=0\)

OpenStudy (ksaimouli):

i think ur correct

OpenStudy (anonymous):

that's what i got (besides the fact that i missed the one lol)

OpenStudy (anonymous):

so then i made my first function over 2squrtxy

OpenStudy (anonymous):

just by multiplyinh it by \[\sqrt{xy}\]

OpenStudy (anonymous):

i just need to know how to further simplify

OpenStudy (ksaimouli):

no take derivative for that one too

OpenStudy (ash2326):

Yeah, the third step is correct. Just find y' now and it's done

OpenStudy (anonymous):

and i crossed multiplied so now i have no denominators to just have: \[(x+y)^\frac{ -1 }{ 2 }*(1+y')*\sqrt{xy}+y+y'+xy'=2\sqrt{xy}\]

hartnn (hartnn):

\(\frac{ (x+y)^\frac{ -1 }{ 2 }*(1+y')*\sqrt{xy} }{ 2\sqrt{xy} }+\frac{ y }{ 2\sqrt{xy} }+\frac{ xy' }{ 2\sqrt{xy} }=0 \\ \large \implies \frac{(x+y)^{-1/2}(1+y')\sqrt{xy}+y+xy'}{2\sqrt{xy}}=0\)

hartnn (hartnn):

then just do numerator =0 and isolate y'

OpenStudy (anonymous):

thats what i had, so then i cross multiplied to get what i just wrote above.

hartnn (hartnn):

isn't right side = 0 ?

OpenStudy (anonymous):

i can't factor out y' though, so do i just divide by everything?

OpenStudy (anonymous):

well i cross multiplied to get rid of my denominator.

OpenStudy (anonymous):

not sure what to do next!

hartnn (hartnn):

\(\frac{ (x+y)^\frac{ -1 }{ 2 }*(1+y')*\sqrt{xy} }{ 2\sqrt{xy} }+\frac{ y }{ 2\sqrt{xy} }+\frac{ xy' }{ 2\sqrt{xy} }=0 \\ \large \implies \frac{(x+y)^{-1/2}(1+y')\sqrt{xy}+y+xy'}{2\sqrt{xy}}=0 \\ \huge \implies (x+y)^{-1/2}(1+y')\sqrt{xy}+y+xy'=0 \)

hartnn (hartnn):

\(\large \implies (x+y)^{-1/2}(1+y')\sqrt{xy}+y+xy'=0\)

OpenStudy (anonymous):

why did you get rid of your denominator?

hartnn (hartnn):

see, right side =0 a/b=0 a=0

OpenStudy (anonymous):

that confuses me that i just drop the denominator ... and i still can't isolate y' :/